OpenStudy (anonymous):

Having trouble with an integral:

OpenStudy (amistre64):

integrals can be troublesome

OpenStudy (anonymous):

\[\int\limits_{ }^{} \frac{1}{P(P-5)}dP\]

OpenStudy (amistre64):

decomp

OpenStudy (anonymous):

i split it to be 1/p and 1/(p-5) and tried to integrate by parts but it doesn't seem to work out nicely. any help would be appreciated

mathslover (mathslover):

\[\large{\int {\frac{1}{P(P-5)}dP}}\] \[\large{\int {\frac{1}{P}dP}\int \frac{1}{P-5}dP}\]

OpenStudy (anonymous):

Whoa Nelly

mathslover (mathslover):

answer will be zero?

OpenStudy (anonymous):

you can't split an integral like that

OpenStudy (amistre64):

\[\frac{1}{P(P-5)}=\frac{A}{P}+\frac{B}{(P-5)}\] \[1=A(P-5)+BP\] when P=0, A=-1/5 when P=5; B=1/5

OpenStudy (anonymous):

i just realised that its probably partial fractions

OpenStudy (anonymous):

well there you go :)

OpenStudy (amistre64):

\[-\frac15\int\frac1Pdp~+~\frac15\int\frac1{P-5}dp\]